# Hi-EF: Benchmarking Emotion Forecasting in Human-interaction

### v1

@@ 5.1 Experiment Setup
5.2 Experimental Results
Experiment Results of EF task. We present the experimental results of the EF task in Table 2. Our primary aim in these experiments is to evaluate the impact of different vision encoders and various intra-video and inter-video fusion strategies on the accuracy of EF. Firstly, we analyze the effect of vision encoders on the experimental results. Since ResNet18 and I3D as vision encoders cannot be aligned with the other two modalities, we only analyze the results using the video modality. We did not employ particularly complex fusion strategies for these two vision encoders, and they did not achieve superior performance. Next, we focus on the results using ViT as the vision encoder. First, we analyze the modalities. The results indicate that among the three modalities, the video modality plays the most significant role in emotion analysis. However, when the text modality is added, there is a decrease in our method’s performance. We believe this is because, when incorporating both video and text information, the text is treated as noise. However, when audio is also included, the model better understands the text information. We found that adding the text modality improves performance when there is an audio modality with video. When using all three modalities, the WAR improved by 3.82% and the UAR by 1.33% compared to using only the video modality. This indicates that we chose the right three modalities for the best performance in the EF task.
Effect of Intra-Video Fusion. During two-party interactions, emotional influence is not solely affected by facial expressions but also by body posture and the current environment. Consequently, we compared using only facial information with a combination of face, body, and scene information. The results show that the prediction accuracy using the combined face + body + scene information is higher than using facial information alone. This indicates that emotional changes during interactions are indeed influenced by these factors. Furthermore, we employed different fusion strategies to integrate these three types of information. The experimental results demonstrate that using a transformer significantly outperforms the averaging method. This is because the emotional content of the three types of information varies, requiring focused learning to capture the most relevant emotional cues.
Effect of Inter-Video Fusion. Since our MCIS includes multiple layers of information, employing appropriate fusion strategies to learn from these different layers is crucial. We used three fusion strategies: average, Rransformer, and LSTM+Transformer. The results show that the LSTM+Transformer strategy performs the best. This is because LSTM captures the temporal information between the three videos, while the transformer captures the key information that influences emotional changes across the videos. The combination of these two methods allows us to predict the other party’s potential emotions most effectively.
Effect of MCIS. We conducted ablation experiments on MCIS to compare the impact of different inputs on the EF task results which are presented in Table 3. When the input video consists of a single clip, the results show that using clip III yields the best performance. When using two video clips, the combination of the clips II and III produces the best results. This indicates that more recent information is more beneficial for the EF task. When we input three video clips, the results significantly outperform the combination of clips II and III, with a WAR improvement of 2.64% and a UAR improvement of 2.33%. This also demonstrates the effectiveness of our MCIS design.
Derived tasks of Hi-EF: In addition to the EF task, the Hi-EF dataset can also be applied to other affective computing tasks such as DFER, MER, Multilabel-classification and ERC tasks. We have included the experimental results and analyses in the Appendix.
[…]
Input Video Intra Fusion Text Audion Inter Fusion WAR UAR
@@ 5.2 Experimental Results
6 Conclusions and Discussion
In this paper, we have transformed the task of Affective Forecasting from a psychological and economic problem into a Deep Learning challenge. We define a new task called Emotion Forecasting, which focuses on predicting emotions during two-party interactions. To facilitate this task, we propose the Hi-EF dataset and a baseline. The dataset features a novel data format, MCIS, which consists of multilayered-contextual interactive information between two parties. We provide rich emotion-related annotations and three modalities for each MCIS. We believe the EF task holds immense potential.
@@ 6 Conclusions and Discussion
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Abstract.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [No] According to our understanding, Emotion Forecasting (EF) does not appear to lead to any significant societal issues.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] The research conforms with the NeurIPS Code of Ethics.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] No theoretical results.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] No theoretical results.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Code is provided HERE.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We provide the implementation details in Section 5
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] Statistical significance bars are not reported because it would be too computationally expensive.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We provide the implementation details in Section 5
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We cite them.
    2. (b)
      Did you mention the license of the assets? [Yes] We mentioned.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] Dataset is provided HERE .
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] The videos we download and scrape from the internet are permitted to be downloaded and scraped.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] We provide the details of instructions in the supplementary materials.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [No] Crowdsourcing only annotate the video and there may be no risk.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [No] This cannot be disclosed publicly.

### v2

@@ Introduction
Table 2: The comparison of EF experiments result($\%$) with different modalities, backbones, and intra and inter fusion strategies. Bold indicates the best performance, while underlining indicates the second-best performance.
Input Video Intra Fusion Text Audio Inter Fusion WAR UAR
@@
Table 3: Different input strategies of EF experiments results($\%$) on Hi-EF. Bold indicates the best performance, while underlining indicates the second-best performance.
Experimental Results
Experimental Results of EF Task. We present the experimental results of the EF task in Table 2. Our main goal is to study the impact of different vision encoders and intra-/inter-video fusion strategies on EF performance. We first compare the vision encoders. Since ResNet18 and I3D cannot be aligned with the other two modalities, we only report video-only results for them, using simple fusion strategies; consequently, they do not achieve competitive performance. We then focus on ViT-based models and analyze the effect of modalities. The results show that the video modality contributes most to emotion analysis. Adding text alone degrades performance, as the text is treated as noise, whereas adding audio enables the model to better interpret the textual information. Using all three modalities improves WAR by 3.82% and UAR by 1.33% over the video-only setting, indicating that our modality choice is appropriate for the EF task.
Effect of Intra-Video Fusion. During two-party interactions, emotional influence arises not only from facial expressions but also from body posture and the surrounding scene. We therefore compare using only facial information with using the combination of face, body, and scene. The combined setting yields higher prediction accuracy, confirming that these factors jointly affect emotional changes. We also study different fusion strategies for these three types of information. The results show that transformer-based fusion clearly outperforms simple averaging, as it can focus on the most informative emotional cues.
Effect of Inter-Video Fusion. Since MCIS contains multiple layers of information, an appropriate inter-video fusion strategy is crucial. We compare three strategies: average, Transformer, and LSTM+Transformer. LSTM+Transformer performs best, where LSTM captures temporal dependencies among the three clips and the transformer highlights key signals that drive emotional changes. Their combination leads to the most accurate prediction of the other party’s potential emotions.
Effect of MCIS. We conduct ablation experiments on MCIS to study the impact of different input configurations, as reported in Table 3. With a single input clip, clip III achieves the best results. With two clips, the combination of clips II and III performs best, indicating that more recent information is more beneficial for EF. When three clips are used, the performance surpasses that of the II+III setting, with WAR and UAR improvements of 2.64% and 2.33%, respectively, demonstrating the effectiveness of the MCIS design.
Derived Tasks of Hi-EF. Beyond EF, the Hi-EF dataset can also support other affective computing tasks, such as DFER, MER, and ERC. The corresponding experimental results are reported in the Appendix.
Conclusions and Discussion
In this paper, we have transformed the task of Affective Forecasting from a psychological and economic problem into a deep learning challenge. We define a new task called Emotion Forecasting, which focuses on forecasting emotions during two-party interactions. To facilitate this task, we propose the Hi-EF dataset and a baseline. The dataset features a novel data format, MCIS, which consists of multilayered-contextual interactive information between two parties. We provide rich emotion-related annotations and three modalities for each MCIS. Our primary contribution lies in designing the EF task and creating a suitable dataset and data format to accommodate this task, ultimately demonstrating its feasibility. We believe that the EF task and the Hi-EF dataset will provide a new perspective for future research in affective computing.
@@
Acknowledgements
This work was supported by National Natural Science Foundation of China (No.62576109, 62072112), National Key Research and Development Program of China (2023YFC3604802) and National Natural Science Foundation of China (No. 62406075)